A has four keys in his pocket, he tries to unlock a door and tries with each key until he finds the right one, let X be random variable number of keys tried including the right key to open the door,...


a) is x a discrete or continuous random variable? Explain


b) determine probability function for random variable x.


C) what is the probability that the man will try at most 2 keys before he finds the right one?


D) determine the mean and standard deviation of x.


1
Expert's answer
2021-12-05T18:28:43-0500

a)

A random variable "X" has possible values "1,2,3,4" which constitute a finite set. Therefore "x" is a discrete random variable.


b)

If unsuccessful keys are eliminated then it can take at most 4 attempts to open the door with the following probabilities:

Probability that the first key will open the door is

"P(X=1)=1\/4"

Probability that the second key will open the door is

"P(X=2)=(1-1\/4)(1\/3)=1\/4"

Probability that the third key will open the door is

"P(X=3)=(1-1\/4)(1-1\/3)(1\/2)=1\/4"

Probability that the fourth key will open the door is

"P(X=4)=(1-1\/4)(1-1\/3)(1-1\/2)(1)=1\/4"

"X" has the uniform distribution with "p=1\/4"


"P(X=x)=1\/4, x=1,2,3,4"

c)


"P(X\leq 2)=P(X=1)+P(X=2)"

"=1\/4+1\/4=1\/2"

d)


"mean=E(X)=1(\dfrac{1}{4})+2(\dfrac{1}{4})+3(\dfrac{1}{4})+4(\dfrac{1}{4})=\dfrac{5}{2}"

"E(X^2)=1^2(\dfrac{1}{4})+2^2(\dfrac{1}{4})+3^2(\dfrac{1}{4})+4^2(\dfrac{1}{4})=\dfrac{15}{2}"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2"

"=\dfrac{15}{2}-(\dfrac{5}{2})^2=\dfrac{5}{4}"

"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{5}{4}}=\dfrac{\sqrt{5}}{2}"